Write down the equation of the plane containing the point (2 , 3 , 4 ) and parallel to

the vectors

u= (2 , -3 ,2) and v= (0 , 1 , 2) .


1
Expert's answer
2021-10-05T13:47:18-0400
"u\times v=\begin{vmatrix}
 i & j & k \\
 2 & -3 & 2 \\
 0 & 1 & 2 \\
\end{vmatrix}"

"=i\begin{vmatrix}
 -3 & 2 \\
 1 & 2
\end{vmatrix}-j\begin{vmatrix}
 2 & 2 \\
 0 & 2
\end{vmatrix}+k\begin{vmatrix}
 2 & -3 \\
 0 & 1
\end{vmatrix}"

"=-8i-4j+2k"

"n=(-8,-4,2)"

The equation of the plane containing the point (2 , 3 , 4 )


"-8(x-2)-4(y-3)+2(z-4)=0"

"4(x-2)+2(y-3)-(z-4)=0"

"4x-8+2y-6-z+4=0"

"4x+2y-z-10=0"